# Discrete Time Crystal Phase as a Resource for Quantum Enhanced Sensing

### v2

@@ Abstract
Discrete time crystals are a special phase of matter in which time translational symmetry is broken through a periodic driving pulse. Here, we first propose and characterize an effective mechanism to generate a stable discrete time crystal phase in a disorder-free many-body system with indefinite persistent oscillations even in finite-size systems. Then we explore the sensing capability of this system to measure the spin exchange coupling. The results show strong quantum-enhanced sensitivity throughout the time crystal phase. As the spin exchange coupling varies, the system goes through a sharp phase transition and enters a non-time crystal phase in which the performance of the probe considerably decreases. We characterize this phase transition as a second-order type and determine its critical properties through a comprehensive finite-size scaling analysis. The performance of our probe is independent of the initial states and may even benefit from imperfections in the driving pulse.
@@
Introduction.– Symmetry breaking is a fundamental process that shapes our universe, from its early evolution and the formation of elementary particles to various forms of phase transitions in our daily lives. Breaking continuous spatial translation symmetry into a discrete one results in ordinary crystals, where atoms sit in a regular order. In a seminal work by Wilczek Wilczek (2012), the idea of breaking continuous time translation symmetry and the formation of time crystals was proposed. While this proposal proved impossible in equilibrium states of time-independent systems with two-body interactions Bruno (2013); Watanabe and Oshikawa (2015); Kozin and Kyriienko (2019), the spontaneous emergence of a new periodic motion turned out to be possible in periodically driven systems Sacha (2015); Khemani et al. (2016); Else et al. (2016). Breaking discrete time translational symmetry (DTTS) in such systems and forming so-called discrete time crystals (DTC) has become the subject of intensive theoretical Sacha (2015); Khemani et al. (2016); Else et al. (2016); Yao et al. (2017); Russomanno et al. (2017); Ho et al. (2017); Huang et al. (2018); Matus and Sacha (2019); Kshetrimayum et al. (2020); Estarellas et al. (2020); Maskara et al. (2021); Wang et al. (2021); Pizzi et al. (2021); Collura et al. (2022); Huang et al. (2022); Bull et al. (2022); Deng and Yang (2023); Liu et al. (2023); Huang (2023); Giergiel et al. (2023) and experimental Zhang et al. (2017); Choi et al. (2017); Pal et al. (2018); Rovny et al. (2018); Smits et al. (2018); Randall et al. (2021); Keßler et al. (2021); Xu et al. (2021); Kyprianidis et al. (2021); Taheri et al. (2022); Mi et al. (2022); Frey and Rachel (2022); Bao et al. (2024); Shinjo et al. (2024); Liu et al. (2024a, b) research (for reviews see Sacha and Zakrzewski (2017); Else et al. (2020); Khemani et al. (2019); Sacha (2020); Hannaford and Sacha (2022); Zaletel et al. (2023)). In periodically driven systems with a period $T$, DTCs do not correspond to equilibrium states but reveal temporal order where: (i) physical observables evolve with period $gT$ with integer $g{>}1$; (ii) the dynamics are robust against small imperfections in the driving pulse; and (iii) the oscillating behavior persists indefinitely in the thermodynamic limit. The existence of the DTC relies on mechanisms that prohibit the system from absorbing energy from the driving pulse, such as self-trapping, the presence of disorder, gradient magnetic fields, all-to-all or long-range interactions, domain-wall confinement, and quantum scars Sacha and Zakrzewski (2017); Else et al. (2020); Khemani et al. (2019); Sacha (2020); Zaletel et al. (2023); Liu et al. (2023); Kshetrimayum et al. (2020); Russomanno et al. (2017); Pizzi et al. (2021); Collura et al. (2022); Maskara et al. (2021); Deng and Yang (2023); Bull et al. (2022); Huang et al. (2022); Huang (2023). While major proposals focus on the formation and detection of DTCs, the potential application of this phase of matter is yet to be explored. So far, time crystals have been used for simulating complex systems Estarellas et al. (2020), topologically protected quantum computation Bomantara and Gong (2018), designing quantum engines Carollo et al. (2020), metrology in fully-connected graphs Lyu et al. (2020), measuring AC fields Iemini et al. (2023), and system-environment coupling Montenegro et al. (2023).
@@
Strongly correlated many-body systems have been identified as excellent quantum sensors. In particular, various forms of quantum criticality have been used for achieving quantum-enhanced sensitivity beyond the capacity of classical sensors. This includes first-order Raghunandan et al. (2018); Heugel et al. (2019); Yang and Jacob (2019), second-order Zanardi and Paunković (2006); Zanardi et al. (2007); Gu et al. (2008); Zanardi et al. (2008); Invernizzi et al. (2008); Gu (2010); Gammelmark and Mølmer (2011); Skotiniotis et al. (2015); Rams et al. (2018); Wei (2019); Chu et al. (2021); Liu et al. (2021); Montenegro et al. (2021); Mirkhalaf et al. (2021); Di Candia et al. (2021); Salvia et al. (2023), dissipative Fernández-Lorenzo and Porras (2017); Baumann et al. (2010); Baden et al. (2014); Klinder et al. (2015); Rodriguez et al. (2017); Fitzpatrick et al. (2017); Fink et al. (2017); Ilias et al. (2022, 2024); Alipour et al. (2014), topological Budich and Bergholtz (2020); Sarkar et al. (2022); Koch and Budich (2022); Yu et al. (2024), Floquet Mishra and Bayat (2021, 2022), and Stark He et al. (2023); Yousefjani et al. (2023a, 2024) phase transitions. However, the benefits of using criticality-based probes are limited by three major factors: (i) the region over which quantum-enhanced precision is achievable is very narrow; (ii) state preparation, e.g. ground state, near the critical point may require a complex time-consuming procedure; and (iii) the presence of imperfection deteriorates the performance of the sensor. Therefore, any sensing protocol that operates optimally over a reasonably wide region without requiring complex state preparation and being stable against unwanted imperfections is highly desired.
@@
Here, by exploiting the state-of-the-art numerical simulations, we put forward a mechanism for establishing a stable DTC with period-doubling oscillations that persist indefinitely even in finite size systems. While the DTC shows strong robustness to a certain value of imperfection in the pulse, it goes through a sharp second-order phase transition as the spin exchange coupling varies. Relying on this transition, we devise a DTC quantum sensor that benefits from multiple features. First, the probe shows extreme sensitivity to the exchange coupling across the whole DTC phase, achieving quantum-enhanced sensitivity. Second, the probe performance is independent of the initial state. Third, the precision enhances by increasing imperfection in the pulse to the certain value. In addition, we also characterize the non-DTC phase observing features of ergodic phase in the thermodynamic limit.
@@
Quantum parameter estimation.– We begin by recapitulating the theory of quantum parameter estimation that aims to infer an unknown parameter $\omega$ in a Hamiltonian of a probe by observing the evolution of the probe’s state $\rho(\omega)$. The uncertainty in estimating $\omega$, quantified through the standard deviation $\delta\omega$, is lower bounded by quantum Cramér-Rao inequality $\delta\omega{\geq}1/\sqrt{\mathcal{F}_{Q}(\omega)}$ wherein $\mathcal{F}_{Q}(\omega)$ is the quantum Fisher information (QFI). For pure states $\rho(\omega){=}|\psi(\omega)\rangle\langle\psi(\omega)|$ the QFI is given by $\mathcal{F}_{Q}(\omega){=}4\big{(}\langle\partial_{\omega}\psi(\omega)|\partial_{\omega}\psi(\omega)\rangle{-}|\langle\partial_{\omega}\psi(\omega)|\psi(\omega)\rangle|^{2}\big{)}$ Fisher (1922). In classical sensors Fisher information, at best, scales linearly with system size $L$. Exploiting quantum features in sensing the coupling of a $k$-body interacting system allows precision enhancement to $\mathcal{F}_{Q}{\sim}L^{2k}$, known as ultimate precision Boixo et al. (2007).
@@
The model.– We consider a one-dimensional chain that contains $L$ spin-$1/2$ with Ising-type interaction, governed by the following Hamiltonian
@@
Here $J$ is the spin exchange coupling, and $\sigma^{x,y,z}_{j}$ are the Pauli operators. The gradient $zz$ interaction in $H_{I}$ causes off-resonant energy splitting at each site and, therefore, leads the particle’s wave function to localize, reminiscent of the localization which is usually induced by applying a gradient magnetic field Schulz et al. (2019); Morong et al. (2021); He et al. (2023); Yousefjani et al. (2023a, 2024). This localization, characterized by the existence of an extensive number of conserved quantities Alet and Laflorencie (2018); Luitz et al. (2015); Yousefjani and Bayat (2023), is essential to prevent our system from absorbing the energy of the periodic drives Lenarčič et al. (2020); Yousefjani et al. (2023b). In the absence of the localization, any local physical observable becomes featureless, and the system thermalizes D’Alessio and Rigol (2014); Lazarides et al. (2014).
Since $H_{P}$ acts in period $T$, the evolution of the system is described by the Floquet theorem. The Floquet unitary operator for one period evolution is
@@
here $\Omega{=}JT$, and $\Phi$ is tuned to be $\Phi{=}(1{-}\varepsilon)\frac{\pi}{2}$, with $\varepsilon$ as deviation from a $\pi/2$ $x$-rotation. In the following, we show how two main parameters, namely $\Omega$ and $\varepsilon$ play roles in establishing a stable DTC. First, we analytically show that setting $\Omega{=}\pi/2$ results in a stable period doubling DTC that is robust against arbitrary imperfection $\varepsilon$ in the rotating pulse. Then, through comprehensive numerical simulations, we show that as $\Omega$ varies from $\pi/2$, the system goes through a sharp phase transition from a stable DTC to a regime in which DTC order is lost. We explore the possibility of this phase transition to act as resource for quantum sensing.
@@
Discrete Time Crystal.– We begin by highlighting some key features of $H_{I}$. First, $H_{I}$ is diagonalized in the computational basis, namely $H_{I}{=}\sum_{\mathbf{z}=1}^{2^{L}}E_{\mathbf{z}}|\mathbf{z}\rangle\langle\mathbf{z}|$ in which $\{|\mathbf{z}\rangle\}$ represents the $2^{L}$ elements of the computational basis. Any state of the computational basis can be written as $|\mathbf{z}\rangle{=}(\sigma_{1}^{x})^{j_{1}}(\sigma_{2}^{x})^{j_{2}}\cdots(\sigma_{L}^{x})^{j_{L}}|{\uparrow,\uparrow,\cdots,\uparrow}\rangle$ with $\mathbf{z}{=}(j_{1},j_{2},\cdots,j_{L})_{2}$ being the binary representation of the integer $\mathbf{z}$. Second, $[H_{I},\Pi_{j}\sigma_{j}^{x}]{=}0$ which implies that $E_{\mathbf{z}}{=}E_{2^{L}-1-\mathbf{z}}$. Third, for an even number of spins which is considered here, all the eigenvalues $E_{\mathbf{z}}$ are integer numbers that are even (odd) if $L{/2}$ is an even (odd) number. Since $[H_{I},\Pi_{j}\sigma_{j}^{x}]{=}0$, for $\varepsilon{=}0$ which results in $e^{-iH_{P}}{=}\Pi_{j}\sigma_{j}^{x}$, one has a trivial period doubling DTC as $U_{F}^{2}(\Omega{=}\pi/2,\varepsilon{=}0){=}e^{-2iH_{P}}e^{-2i\Omega H_{I}}{=}(-1)^{H_{I}}{=}{\pm}\mathbb{I}$. Consequently, one observes persistent oscillations in typical observables with spontaneously breaking DTTS. For $\varepsilon{\neq}0$, one gets $[H_{I},e^{-iH_{P}}]{\neq}0$. In this case, the reduction of $U_{F}^{2}(\Omega{=}\pi/2,\varepsilon{\neq}0)$ to the identity is not obvious. To study this nontrivial DTC, we focus on the evolution of computational basis over $n$ period cycles and its revival fidelity $F(nT){=}|\langle\mathbf{z}|U_{F}^{n}(\Omega{=}\pi/2,\varepsilon{\neq}0)|\mathbf{z}\rangle|^{2}$ in the system with even number of spins. For a typical computational basis state $|\mathbf{z}\rangle$ with $u_{\mathbf{z}}$ spins down, the free evolution of the system governed by $H_{I}$ imposes a dynamical phase as $e^{-i\Omega H_{I}}|\mathbf{z}\rangle{=}(-i)^{L/2}(-1)^{u_{\mathbf{z}}}|\mathbf{z}\rangle$. Then the first rotating pulse evolves $|\mathbf{z}\rangle$ to a superposition of all the $2^{L}$ elements, each with coefficient $(-i\sin\Phi)^{f}(\cos\Phi)^{L-f}$ wherein $f$ is the number of the flipped spins. After the second period of the evolution, one can show that $\langle\mathbf{z}|U_{F}^{2}(\Omega{=}\pi/2,\varepsilon{\neq}0)|\mathbf{z}\rangle$ is equal with the summation of $2^{L}$ choices of flipping $L$ spins with coefficient $(i)^{L}(-i\sin\Phi)^{2f}(\cos\Phi)^{2(L-f)}$. A straightforward simplification results in $\langle\mathbf{z}|U_{F}^{2}(\Omega{=}\pi/2,\varepsilon{\neq}0)|\mathbf{z}\rangle{=}(-i)^{L}$ and, hence, $F(2T){=}1$, see Supplementary Materials (SM). This calculation shows that regardless of the imperfections in the driving pulse, as long as $\Omega{=}\pi/2$ any initial state returns to itself after time $2T$, therefore, period-doubling oscillations of $F(nT)$ persists indefinitely even in finite size systems.
@@
However, establishing a stable DTC must be independent of fine-tuned Hamiltonian parameters. This obliges us to analyze the effect of a deviation as $\omega{=}|\pi/2{-}\Omega|$. Surprisingly, our comprehensive numerical simulations show that as $\omega$ increases, our system goes through a sharp phase transition from a stable DTC to a region with no spontaneous breaking of DTTS in Eq.(1). Before presenting the main results, some general points related to methodologies need to be clarified. Throughout this Letter, we used the exact diagonalization (ED) computational method for the system of size $L{=}12$ and time-dependent variational principle (TDVP) techniques for finite matrix product state (MPS), using PYTHON package TeNPy Hauschild and Pollmann (2018), for systems of size $L{>}12$. The results are presented for the initial state $|\psi_{0}\rangle{=}|\mathbf{0}\rangle{=}|{\uparrow,\cdots,\uparrow\rangle}$, although, the results are generic and remain valid for other computational basis states too (see SM). In Fig. 1(a), we plot stroboscopic dynamics of the revival fidelity $F(2nT)$ as a function of $n$ for different values of $\omega$ in the system of size $L{=}12$ under a driving pulse with an imperfection of magnitude $\varepsilon{=}0.01$. In the stable DTC phase, happening in the range $\omega{\leq}10^{-2}\pi/2$, one observes $F(2nT){=}1$. For larger values of the deviation, such as $\omega{\geq}0.05\pi/2$, revival fidelity shows nontrivial oscillations, signaling the entrance to a non-DTC region. We characterize this region later. This distinctive behavior with respect to $\omega$ reflects itself in all stroboscopic times as has been depicted in Fig. 1(b). In this panel, we plot the revival fidelity at different stroboscopic times $n{\in}\{2,10,\cdots,100\}$, in a chain of length $L{=}30$ and $\varepsilon{=}0.01$. The inset represents the average fidelity $\overline{F}(nT){=}(1/N)\sum_{n=1}^{N}F(2nT)$ for the considered stroboscopic times. As is obvious from Fig. 1(b), the phase transition between stable DTC and non-DTC region occurs at a specific value of $\omega{=}\omega_{\max}$, dashed line, in all the stroboscopic times. In the following, we first analyze the capability of this phase transition as a resource for quantum sensing. Then, we complete this analysis by extracting the critical features of the quantum phase transition using a well-established mechanism that identifies the type of transition as a second-order one.

### v3

@@ Abstract
Discrete time crystals are a special phase of matter in which time translational symmetry is broken through a periodic driving pulse. Here, we first propose and characterize an effective mechanism to generate a stable discrete time crystal phase in a disorder-free many-body system with indefinite persistent oscillations even in finite-size systems. Then we explore the sensing capability of this system to measure the spin exchange coupling. The results show strong quantum-enhanced sensitivity throughout the time crystal phase. As the spin exchange coupling varies, the system goes through a sharp phase transition and enters a non-time crystal phase in which the performance of the probe considerably decreases. We characterize this phase transition as a second-order type and determine its critical properties through a comprehensive finite-size scaling analysis. The performance is independent of the initial states and may even benefit from imperfections in the driving pulse. A simple set of projective measurements can capture the quantum-enhanced sensitivity.
1 Introduction
@@
Symmetry breaking is a fundamental process that shapes our universe, from its early evolution and the formation of elementary particles to various forms of phase transitions in our daily lives. Breaking continuous spatial translation symmetry into a discrete one results in ordinary crystals, where atoms sit in a regular order. In a seminal work by Wilczek [1], the idea of breaking continuous time translation symmetry and the formation of time crystals was proposed. While this proposal proved impossible in equilibrium states of time-independent systems with two-body interactions [2, 3, 4], the spontaneous emergence of a new periodic motion turned out to be possible in periodically driven systems [5, 6, 7]. Breaking discrete time translational symmetry (DTTS) in such systems and forming so-called discrete time crystals (DTC) has become the subject of intensive theoretical [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24] and experimental [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40] research (for reviews see [41, 42, 43, 44, 45, 46]). In periodically driven systems with a period $T$, DTCs do not correspond to equilibrium states but reveal temporal order where: (i) physical observables evolve with period $gT$ with integer $g{>}1$; (ii) the dynamics are robust against small imperfections in the driving pulse; and (iii) the oscillating behavior persists indefinitely in the thermodynamic limit. The existence of the DTC relies on mechanisms that prohibit the system from absorbing energy from the driving pulse, such as self-trapping, the presence of disorder, gradient magnetic fields, all-to-all or long-range interactions, domain-wall confinement, and quantum scars [41, 42, 43, 44, 46, 22, 13, 9, 17, 18, 15, 21, 20, 19, 23]. The study has been also extended to non-Hermitian physics [47]. While major proposals focus on the formation and detection of DTCs, the potential application of this phase of matter is yet to be explored. So far, time crystals have been used for simulating complex systems [14], topologically protected quantum computation [48], designing quantum engines [49], metrology in fully-connected graphs [50], measuring AC fields [51], and system-environment coupling [52, 53].
@@
Strongly correlated many-body systems have been identified as excellent quantum sensors. In particular, various forms of quantum criticality have been used for achieving quantum-enhanced sensitivity beyond the capacity of classical sensors. This includes first-order [54, 55, 56], second-order [57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72], dissipative [73, 74, 75, 76, 77, 78, 79, 80, 81, 82], topological [83, 84, 85, 86], Floquet [87, 88], and Stark [89, 90, 91] phase transitions. For a reference see [92]. However, the benefits of using criticality-based probes are limited by three major factors: (i) the region over which quantum-enhanced precision is achievable is very narrow; (ii) state preparation, e.g. ground state, near the critical point may require a complex time-consuming procedure; and (iii) the presence of imperfection deteriorates the performance of the sensor. Therefore, any sensing protocol that operates optimally over a reasonably wide region without requiring complex state preparation and being stable against unwanted imperfections is highly desired.
@@
Here, by exploiting the state-of-the-art numerical simulations, we put forward a mechanism for establishing a stable DTC with period-doubling oscillations that persist indefinitely even in finite size systems. While the DTC shows strong robustness to a certain value of imperfection in the pulse, it goes through a sharp second-order phase transition as the spin exchange coupling varies. Relying on this transition, we devise a DTC quantum sensor that benefits from multiple features. First, the probe shows extreme sensitivity to the exchange coupling across the whole DTC phase, resulting quantum-enhanced sensitivity. Second, the probe performance is independent of the initial state. Third, the precision enhances by increasing imperfection in the pulse to a certain value. Forth, a simple set of projective measurements allows to capture the quantum-enhanced sensitivity. In addition, we also characterize the non-DTC phase observing features of ergodic phase in the thermodynamic limit.
2 Quantum parameter estimation
@@
We begin by recapitulating the theory of quantum parameter estimation that aims to infer an unknown parameter $\omega$ in a Hamiltonian of a probe by observing the evolution of the probe’s state $\rho(\omega)$. The uncertainty in estimating $\omega$, quantified through the standard deviation $\delta\omega$, is lower bounded by Cramér-Rao inequality $\delta\omega{\geq}1/\sqrt{\mathcal{F}_{C}(\omega)}$ wherein $\mathcal{F}_{C}(\omega){=}\sum_{r}p_{r}(\omega)[\partial_{\omega}\ln p_{r}(\omega)]^{2}$ is the Classical Fisher information (CFI). Here $p_{r}(\omega){=}{\rm Tr}[\rho(\omega)\Pi_{r}]$ is the outcome probability of measuring the probe using operators $\{\Pi_{r}\}$. Optimizing the measurement operators leads to $\delta\omega{\geq}1/\sqrt{\mathcal{F}_{Q}(\omega)}$, known as the quantum Cramér-Rao inequality wherein $\mathcal{F}_{Q}(\omega)$ is the quantum Fisher information (QFI). For pure states $\rho(\omega){=}|\psi(\omega)\rangle\langle\psi(\omega)|$ the QFI is given by $\mathcal{F}_{Q}(\omega){=}4\big{(}\langle\partial_{\omega}\psi(\omega)|\partial_{\omega}\psi(\omega)\rangle{-}|\langle\partial_{\omega}\psi(\omega)|\psi(\omega)\rangle|^{2}\big{)}$ [93]. In classical sensors Fisher information, at best, scales linearly with system size $L$. Exploiting quantum features in sensing the coupling of a $k$-body interacting system allows precision enhancement to $\mathcal{F}_{Q}{\sim}L^{2k}$, known as ultimate precision [94].
3 The model
@@
We consider a one-dimensional chain that contains $L$ spin-$1/2$ with Ising-type interaction, governed by the following Hamiltonian
@@
Here $J$ is the spin exchange coupling, and $\sigma^{x,y,z}_{j}$ are the Pauli operators. The gradient $zz$ interaction in $H_{I}$ causes off-resonant energy splitting at each site and, therefore, leads the particle’s wave function to localize, reminiscent of the localization which is usually induced by applying a gradient magnetic field [95, 96, 89, 90, 91]. This localization, characterized by the existence of an extensive number of conserved quantities [97, 98, 99], is essential to prevent our system from absorbing the energy of the periodic drives [100, 101]. In the absence of the localization, any local physical observable becomes featureless, and the system thermalizes [102, 103]. Since $H_{P}$ acts in period $T$, the Floquet unitary operator for one period evolution is
@@
here $\Omega{=}JT$, and $\Phi$ is tuned to be $\Phi{=}(1{-}\varepsilon)\frac{\pi}{2}$, with $\varepsilon$ as deviation from a $\pi/2$ $x$-rotation. In the following, we show how two main parameters, namely $\Omega$ and $\varepsilon$ play roles in establishing a stable DTC. First, we analytically show that setting $\Omega{=}\pi/2$ results in a stable period doubling DTC that is robust against arbitrary imperfection $\varepsilon$. Then, through comprehensive numerical simulations, we show that as $\Omega$ varies from $\pi/2$, the system goes through a sharp phase transition from a stable DTC to a regime in which DTC order is lost. We explore the possibility of this phase transition to act as resource for quantum sensing.
@@ 4 Discrete Time Crystal
We begin by highlighting some key features of $H_{I}$. First, $H_{I}$ is diagonalized in the computational basis $\{|\mathbf{z}\rangle\}$, namely $H_{I}{=}\sum_{\mathbf{z}=1}^{2^{L}}E_{\mathbf{z}}|\mathbf{z}\rangle\langle\mathbf{z}|$. Here $|\mathbf{z}\rangle{=}(\sigma_{1}^{x})^{j_{1}}(\sigma_{2}^{x})^{j_{2}}\cdots(\sigma_{L}^{x})^{j_{L}}|{\uparrow,\uparrow,\cdots,\uparrow}\rangle$ with $\mathbf{z}{=}(j_{1},j_{2},\cdots,j_{L})_{2}$ being the binary representation of the integer $\mathbf{z}$. Second, $[H_{I},\Pi_{j}\sigma_{j}^{x}]{=}0$ which implies that $E_{\mathbf{z}}{=}E_{2^{L}-1-\mathbf{z}}$. Third, for an even number of spins which is considered here, all the eigenvalues $E_{\mathbf{z}}$ are integer numbers that are even (odd) if $L{/2}$ is an even (odd) number. Since $[H_{I},\Pi_{j}\sigma_{j}^{x}]{=}0$, for $\varepsilon{=}0$ which results in $e^{-iH_{P}}{=}\Pi_{j}\sigma_{j}^{x}$, one has a trivial period doubling DTC as $U_{F}^{2}(\Omega{=}\pi/2,\varepsilon{=}0){=}e^{-2iH_{P}}e^{-2i\Omega H_{I}}{=}(-1)^{H_{I}}{=}{\pm}\mathbb{I}$. Consequently, one observes persistent oscillations in typical observables with spontaneously breaking DTTS. For $\varepsilon{\neq}0$, one gets $[H_{I},e^{-iH_{P}}]{\neq}0$. In this case, the reduction of $U_{F}^{2}(\Omega{=}\pi/2,\varepsilon{\neq}0)$ to the identity is not obvious. To study this nontrivial DTC, we focus on dynamic of $|\mathbf{z}\rangle$ over $n$ period cycles and its revival fidelity $F(nT){=}|\langle\mathbf{z}|U_{F}^{n}(\Omega{=}\pi/2,\varepsilon{\neq}0)|\mathbf{z}\rangle|^{2}$. For a typical $|\mathbf{z}\rangle$ with $u_{\mathbf{z}}$ spins down, one has $e^{-i\Omega H_{I}}|\mathbf{z}\rangle{=}(-i)^{L/2}(-1)^{u_{\mathbf{z}}}|\mathbf{z}\rangle$. Then the first rotating pulse evolves $|\mathbf{z}\rangle$ to a superposition of all the $2^{L}$ elements, each with coefficient $(-i\sin\Phi)^{f}(\cos\Phi)^{L-f}$ wherein $f$ is the number of the flipped spins. After the second period of the evolution, one can show that $\langle\mathbf{z}|U_{F}^{2}(\Omega{=}\pi/2,\varepsilon{\neq}0)|\mathbf{z}\rangle$ is the summation of $2^{L}$ choices of flipping $L$ spins with coefficient $(i)^{L}(-i\sin\Phi)^{2f}(\cos\Phi)^{2(L-f)}$. A straightforward simplification results in $\langle\mathbf{z}|U_{F}^{2}(\Omega{=}\pi/2,\varepsilon{\neq}0)|\mathbf{z}\rangle{=}(-i)^{L}$ and, hence, $F(2T){=}1$, see appendix A. This evidences that regardless of the imperfections $\varepsilon$, as long as $\Omega{=}\pi/2$ any initial state returns to itself after $2T$, therefore, period-doubling oscillations of $F(nT)$ persists indefinitely even in finite size systems.
@@
However, establishing a stable DTC must be independent of fine-tuned Hamiltonian parameters. This obliges us to analyze the effect of a deviation as $\omega{=}|\pi/2{-}\Omega|$. Surprisingly, our comprehensive numerical simulations show that as $\omega$ increases, our system goes through a sharp phase transition from a stable DTC to a region with no spontaneous breaking of DTTS in Eq.(1). Before detailing the main findings, some methodological notes must be clarified. We used the exact diagonalization (ED) for $L{=}12$ and time-dependent variational principle (TDVP) techniques for finite matrix product state (MPS), using PYTHON package TeNPy [104], for $L{>}12$. The results are presented for the initial state $|\psi_{0}\rangle{=}|\mathbf{0}\rangle{=}|{\uparrow,\cdots,\uparrow\rangle}$, although, the results are generic and remain valid for other computational basis states too (see appendix B). In Fig. 1(a), we plot stroboscopic dynamics of the revival fidelity $F(2nT)$ as a function of $n$ for various $\omega$ in the system of size $L{=}12$ under a driving pulse with $\varepsilon{=}0.01$. In the stable DTC phase, happening in the range $\omega{\leq}10^{-2}\pi/2$, one observes $F(2nT){=}1$. For larger values of the deviation, such as $\omega{\geq}0.05\pi/2$, revival fidelity shows nontrivial oscillations, signaling the entrance to a non-DTC region. We characterize this region later. This distinctive behavior with respect to $\omega$ reflects itself in all stroboscopic times as has been depicted in Fig. 1(b). In this panel, we plot the revival fidelity at different stroboscopic times $n{\in}\{2,10,\cdots,100\}$, in a chain of length $L{=}30$ and $\varepsilon{=}0.01$. The inset represents the average fidelity $\overline{F}(nT){=}(1/N)\sum_{n=1}^{N}F(2nT)$ for the considered stroboscopic times. As is obvious from Fig. 1(b), the phase transition between stable DTC and non-DTC region occurs at a specific value of $\omega{=}\omega_{\max}$, dashed line, in all the stroboscopic times. In the following, we first analyze the capability of this phase transition as a resource for quantum sensing. Then, we complete this analysis by extracting the critical features of the quantum phase transition using a well-established mechanism that identifies the type of transition as a second-order one.
5 DTC sensor